The length of a rectangle is twice the width. If the perimeter of the rectangle is 30 units, find the area of the garden.

Answer by stanbon(75887)

The length of a rectangle is twice the width. If the perimeter of the rectangle is 30 units, find the area of the garden.
------------------------
L = 2*W
2(L+W) = 30
----------------
L + W = 15
-----
2W + W = 15
3W = 15
Width = 5 units
L = 2*W = 10 units
--------------
Cheers,
Stan H.
------------